Right now the price of Netflix stock is $100, its beta with the market is 2. The expected price next year is $120. The risk free rate is 2% and the market risk premium is 8%. 1. If Winston (the representative investor) believes CAPM correctly describes the risk return tradeoff, will he be willing to buy Netflix stock at the current price? a. i. Yes, Winston will buy exactly the amount suppliede ii. Yes, but Winston will demand more than the amount supplied at $1004 ii. No, Winston will not buy Netflix at $100 v. Not enough information b. If CAPM holds and assuming its beta stays constant, at what price will Winston demand Netflix in an amount exactly equal to supply. (hint: what price today would correctly reflect the CAPM risk?)
Answer #1

(1a)

Risk-Free Rate = 2 %, Market Risk Premium = MRP = 8 % and Beta = 2

Using CAPM, the (minimum) required rate of return = Risk-Free Rate + Beta x MRP = 2 + 2 x 8 = 18 % with this return being the total expected equilibrium return. This is also the minimum return that any investor in Netflix would demand.

Initial Stock Price = $ 100 and Expected Price after 1 year = $ 120

Actual Market Return = (120 - 100) / 100 = 0.2 or 20 %

As the actual market return is greater than the investor's minimum required return (equilibrium total expected return), Wilson will not only buy Netflix at $ 100 but will also demand more than the amount supplied at $ 100, as he is getting reater return (in the form of actual market return) than expected (minimum required return/equilibrium total return). Hence, the correct option is (i).

(1b) Equilibrium Total Return = 18 % and Expected price after 1 year = $ 120

Fair Price of Netflix today (as per equilibrium total return) = 120 / 1.18 = $ 101.695 ~ $ 101.7

This price fairly reflects the CAPM risk. as the price is calculated using the CAPM return rate of 18%.